# Unified Unbiased Variance Estimation for MMD: Robust Finite-Sample Performance with Imbalanced Data and Exact Acceleration under Null and Alternative Hypotheses

### v1

@@ 1 Introduction
The fundamental problem of nonparametric two-sample testing is to detect differences between two probability distributions given access only to samples drawn from them. Maximum Mean Discrepancy (MMD) is a kernel-based nonparametric statistic proposed for this task [6, 11]. When equipped with a characteristic kernel, MMD defines a metric on the space of probability distributions, and thus equals zero if and only if the two distributions are identical [26].
@@
The asymptotic behavior of the unbiased MMD statistic is, however, non-trivial. Under the null hypothesis, the statistic is a degenerate U-statistic, with a limiting distribution given by an infinite weighted sum of chi-square random variables. Under the alternative hypothesis, it converges to a Gaussian distribution [11, 3, 25]. In practice, sample sizes from the two distributions often differ. Existing work has addressed various special cases: Gretton et al. [11] provided the variance under the null hypothesis with equal sample sizes; Sutherland and Deka [27] characterized the unbiased MMD estimator as a U-statistic under balanced sampling; and Wei et al. [29] proposed a method under the alternative hypothesis with unequal sample sizes, treating the MMD estimator as a unified generalized U-statistic [18, 24]. In this context, a unified formulation capable of providing unbiased variance estimation under both hypotheses while accommodating unequal sample sizes remains of significant theoretical and practical interest.
@@
To address these limitations, our approach employs the Hoeffding decomposition to explicitly construct orthogonal projection terms, providing a streamlined representation of the MMD variance estimator. This formulation implicitly incorporates the covariance components present in [29], while simplifying the overall structure. The orthogonal structure further enables a unified and smooth modeling of the variance transition across the null and alternative hypotheses. Beyond its theoretical elegance, the unbiased estimator’s algebraic clarity facilitates a concise, tensorized implementation in modern deep learning frameworks [1, 23], offering a practical and interpretable objective for kernel optimization aimed at maximizing test power [13, 28, 21, 7].
@@
Beyond the theoretical perspective, accurately estimating the MMD variance poses significant computational challenges, as naive summation incurs quadratic complexity, straining both memory and computational resources and limiting the practical deployment of recent theoretical advances [22]. Existing acceleration strategies exhibit clear trade-offs: global approximation methods, such as subsampling or random feature projections, scale well to high-dimensional data but introduce inherent approximation errors [30, 31]. Analytical acceleration offers an alternative, restructuring independent summations into ordered linear recursions [2, 17], a strategy previously applied in classical statistics for measures like Kendall’s tau and distance covariance to reduce complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$ [19, 16]. In the MMD context, Bodenham and Kawahara [5] leveraged the semigroup property of the Laplace kernel to achieve linear-time acceleration of the MMD mean (euMMD) for univariate settings. However, this computational advantage has not yet been extended to unbiased variance estimation, where handling complex second-order terms, including Frobenius norms of kernel matrices, norms of row-sum vectors, and cross-matrix interactions, prevents existing linear-recursion schemes from providing exact results [27].
@@ 2.1 Maximum Mean Discrepancy
Assume $\boldsymbol{X},\boldsymbol{X^{\prime}}{\sim}\mathbb{P}$ and $\boldsymbol{Y},\boldsymbol{Y^{\prime}}{\sim}\mathbb{Q}$ are independent copies. The maximum mean discrepancy is defined as ([11, 22]):
@@
Given i.i.d. finite samples $\mathbf{X}=\left\{x_{i}\right\}_{i=1}^{n}$ and $\mathbf{Y}=\left\{y_{j}\right\}_{j=1}^{m}$ from $\mathbb{P}$ and $\mathbb{Q}$, respectively, an unbiased estimator of the squared MMD is given by ([6]):
@@
which is a linear combination of two single-sample $U$-statistics and one generalized U-statistic ([25]).
The effectiveness of MMD hinges on the use of characteristic kernels, for which the kernel mean embedding uniquely represents probability distributions in the associated reproducing kernel Hilbert space (RKHS) ([26, 9]). Under this condition, MMD induces a proper metric on distributions, ensuring that any difference between them is detectable ([8, 10]).
@@
Finally, although $\widehat{\mathrm{MMD}}^{2}$ enjoys favorable theoretical properties, its computation relies on the explicit construction of kernel matrices, resulting in a computational complexity of $\mathcal{O}(d(n+m)^{2})$. Letting $N=n+m$, this complexity can be written as $\mathcal{O}(dN^{2})$([5]). In large-scale datasets or scenarios where moderate sample sizes require repeated evaluations, such quadratic computational cost severely limits the practical applicability of MMD-based statistical inference.
@@ 2.2 U-statistics and Hoeffding Decomposition
Depending on the sample source, U-statistics can be classified into single-sample U-statistics and generalized U-statistics. Both types form the theoretical foundation of the MMD estimator ([25]).
Let finite samples be $\mathbf{X}=\{x_{i}\}_{i=1}^{n}$ i.i.d. from $\boldsymbol{X}\sim\mathbb{P}$ and $\mathbf{Y}=\{y_{j}\}_{j=1}^{m}$ i.i.d. from $\boldsymbol{Y}\sim\mathbb{Q}$, where $\boldsymbol{X}$ and $\boldsymbol{Y}$ are independent. Consider a function $\theta(\boldsymbol{X})=\mathbb{E}_{\boldsymbol{X}}[h(x_{1},\dots,x_{r})]$, whose unbiased estimator is
@@
where $r$ denotes the order of the U-statistic and $n$ the sample size.
@@
The orthogonal projection components $g_{c}$ are recursively defined by ([14, 4]):
@@
Define $S_{c,n}=\sum_{1\leq i_{1}<\cdots<i_{c}\leq n}g_{c}(x_{i_{1}},\dots,x_{i_{c}})$, then the Hoeffding decomposition yields by ([15]):
@@
where $r_{1}$ and $r_{2}$ denote the orders of the U-statistic for the two samples, respectively.
@@
Define the partial sums as:
$S_{c_{1},c_{2},n,m}=\sum_{1\leq i_{1}<\dots<i_{c_{1}}\leq n}\sum_{1\leq j_{1}<\dots<j_{c_{2}}\leq m}g_{c_{1},c_{2}}(x_{i_{1}},\dots,x_{i_{c_{1}}};y_{j_{1}},\dots,y_{j_{c_{2}}}).$ (12)
The generalized Hoeffding decomposition is then expressed as:
@@ 2.3 Lossless acceleration of the unbiased MMD estimator with the Laplace kernel
Algorithm 1 TriSSL: Triangular Sorted Sum of Lapliacians
@@ 3.1 Variance derivation of the unbiased MMD estimator
According to the Hoeffding decomposition theorem (Eq. 8 and 13), the deviation of each component can be expanded into a sum of linear terms and second-order residual terms.
@@
Unlike the approach in [29], which directly applies the general properties of generalized U-statistics by treating the entire MMD as a single monolithic entity, we leverage the natural form of the MMD estimator as a combination of distinct U-statistic components. By combining Eq. 19, 20, and 21, the deviation of the unbiased estimator can be expressed as:
@@
Since different orders in the Hoeffding decomposition are orthogonal, i.e., $\mathrm{Cov}(H_{1},H_{2})=0$ ([15]), the variance of $\widehat{\mathrm{MMD}}^{2}$ admits the exact decomposition
@@ 3.2.1 Simplification and unbiased estimation of first-order projection terms
where $\widehat{\operatorname{Var}}(\mathcal{U})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$ represent the population variances of the influence functions for distributions $\mathbb{P}$ and $\mathbb{Q}$, respectively. To realize this empirically, we replace the continuous conditional expectations with their discrete sample-based counterparts:
@@
where $\bar{\hat{\mathcal{U}}}=\sum_{i=1}^{n}\hat{\mathcal{U}}(x_{i})\big/n$ and an analogous estimator can be obtained for $\hat{\mathcal{U}}(x_{i})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$:
$\hat{\mathcal{V}}(y_{j})=\frac{1}{m-1}\sum_{j^{\prime}\neq j}k(y_{j},y_{j^{\prime}})-\frac{1}{n}\sum_{i=1}^{m}k(x_{i},y_{j})$ (31)
@@
where we avoid the appearance of covariance terms (which appear in [29]) by computing the variance via continuous projections, thereby simplifying the structure of the unbiased estimator.
@@ 3.2.2 Simplification and unbiased estimation of second-order residual terms
By applying the inclusion-exclusion principle, the unbiased estimator can be expressed entirely in terms of unconstrained matrix operations ([27]). Specifically, using the falling factorial $(n)_{k}$ and the centered kernel matrices $\tilde{\mathbf{K}}_{XX}$ and $\tilde{\mathbf{K}}_{YY}$ (with zero diagonals), the expectation $\mathbb{E}[g_{2}^{2}]$ can be written as a linear combination of the squared Frobenius norms, the squared row and column sums, and the squared total sum of the kernel matrices.
@@ 3.2.3 The unbiased estimator of MMD variance
Using the simplification results provided in Eq. 30324445, and 46, the full analytical expression of $\operatorname{Var}(\widehat{\mathrm{MMD}}^{2})$ is given by
@@
Under the null hypothesis $\mathcal{H}_{0}:\mathbb{P}=\mathbb{Q}$, the unbiased MMD statistic exhibits a pronounced "degenerate" behavior. According to Eq. (3.2.3), its theoretical variance decomposes into a sum of the first-order projection term $\mathcal{T}_{1}$ and the second-order residual term $\mathcal{T}_{2}$. For $\mathcal{T}_{1}$,
@@
$\mathcal{V}\left(y\right)$ $:=\mathbb{E}_{\boldsymbol{Y}^{\prime}}\left[k\left(y,\boldsymbol{Y}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{X}}\left[k\left(\boldsymbol{X},y\right)\right]\equiv 0.$ (49)
Consequently $\operatorname{Var}(\mathcal{T}_{1})\equiv 0$, and the population variance of the first-order term $\mathcal{T}_{1}$ drops from $O(n^{-1})$ to $O(n^{-2})$([11, 29]).
However, it is important to emphasize that this theoretical degeneracy does not imply the disappearance of empirical projection fluctuations. In finite sample observations, we compute the empirical projection $\widehat{\mathcal{U}}(x_{i})$ and $\widehat{\mathcal{V}}(y_{j})$ based on sample averages (Eq. 29 and 31). Due to cross-coupling effects among samples, the empirical projection retains substantial variance. As a result, the estimator in Eq. LABEL:uVarEst is generally larger than the one in [11].
@@ 4.2 Accelerated computation of variance components
4.2.1 Accelerated computation of first-order components.
@@
4.2.2 Accelerated computation of second-order components.
All relevant matrix norms and sums can be expressed directly in terms of these row sums (Eq. 4445, and  46):

### v2

# Unified Unbiased Variance Estimation for Maximum Mean Discrepancy: Robust Finite-Sample Performance with Imbalanced Data and Exact Acceleration under Null and Alternative Hypotheses

@@ 1 Introduction
The fundamental problem of nonparametric two-sample testing is to detect differences between two probability distributions given access only to samples drawn from them. Maximum Mean Discrepancy (MMD) is a kernel-based nonparametric statistic proposed for this task [6, 11]. When equipped with a characteristic kernel, MMD defines a metric on the space of probability distributions, and thus equals zero if and only if the two distributions are identical [22, 27]. It transforms the distance measure between any two distributions into a linear operation in the Hilbert space defined by a kernel function, which greatly facilitates computations. For this reason, this method has attracted increasing attention in recent years. In particular, MMD has been successfully applied to tasks such as domain adaptation[23] and generative modeling[29], and has shown promising potential in other areas, such as Bayesian inference and model calibration tasks [31, 32].
@@
The inferential reliability of MMD, particularly in constructing hypothesis tests and confidence intervals, depends crucially on an accurate characterization of its variance. However, the asymptotic behavior of the unbiased MMD statistic is non-trivial, leading to distinct variance structures under different regimes. Under the null hypothesis, the statistic is a degenerate U-statistic, with a limiting distribution given by an infinite weighted sum of chi-square random variables. Under the alternative hypothesis, it converges to a Gaussian distribution [11, 3]. This dichotomy makes variance estimation challenging, especially in practice where sample sizes from the two distributions often differ. Existing work has addressed various special cases: Gretton et al. [11] provided the unbiased variance estimator under the null hypothesis with equal sample sizes; Sutherland and Deka [28] characterized the unbiased MMD estimator as a U-statistic under balanced sampling; and Wei et al. [30] proposed a method under the alternative hypothesis with unequal sample sizes by treating the MMD estimator as a unified generalized U-statistic [18, 25], although their variance estimator is based on asymptotic approximations and remains biased in finite-sample scenarios. In this context, a unified formulation capable of providing unbiased variance estimation under both hypotheses while accommodating unequal sample sizes remains of significant theoretical and practical interest.
@@
To address these limitations, our approach employs the Hoeffding decomposition to explicitly construct orthogonal projection terms [26, 14], providing a streamlined representation of the MMD variance estimator. This formulation implicitly incorporates the covariance components present in [30], while simplifying the overall structure. The orthogonal structure further enables a unified and smooth modeling of the variance transition across the null and alternative hypotheses. Beyond its theoretical elegance, the unbiased estimator’s algebraic clarity facilitates a concise, tensorized implementation in modern deep learning frameworks [1, 24], offering a practical and interpretable objective for kernel optimization aimed at maximizing test power [29, 13, 21, 7].
@@
Beyond the theoretical perspective, accurately estimating the MMD variance poses significant computational challenges, as naive summation incurs quadratic complexity, straining both memory and computational resources and limiting the practical deployment of recent theoretical advances. Existing acceleration strategies exhibit clear trade-offs: global approximation methods, such as subsampling or random feature projections, scale well to high-dimensional data but introduce inherent approximation errors [33, 34]. Analytical acceleration offers an alternative, restructuring independent summations into ordered linear recursions [2, 17], a strategy previously applied in classical statistics for measures like Kendall’s tau and distance covariance to reduce complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$ [19, 16]. In the MMD context, Bodenham and Kawahara [5] leveraged the semigroup property of the Laplace kernel to achieve linear-time acceleration of the MMD mean (euMMD) for univariate settings. However, this computational advantage has not yet been extended to unbiased variance estimation, where handling complex second-order terms, including Frobenius norms of kernel matrices, norms of row-sum vectors, and cross-matrix interactions, prevents existing linear-recursion schemes from providing exact results [28].
@@ 2.1 Maximum Mean Discrepancy
Assume $\boldsymbol{X},\boldsymbol{X^{\prime}}{\sim}\mathbb{P}$ and $\boldsymbol{Y},\boldsymbol{Y^{\prime}}{\sim}\mathbb{Q}$ are independent copies. The maximum mean discrepancy is defined as [11, 22]:
@@
Given i.i.d. finite samples $\mathbf{X}=\left\{x_{i}\right\}_{i=1}^{n}$ and $\mathbf{Y}=\left\{y_{j}\right\}_{j=1}^{m}$ from $\mathbb{P}$ and $\mathbb{Q}$, respectively, an unbiased estimator of the squared MMD is given by [6]:
@@
which is a linear combination of two single-sample U-statistics and one generalized U-statistic [26].
The effectiveness of MMD hinges on the use of characteristic kernels, for which the kernel mean embedding uniquely represents probability distributions in the associated reproducing kernel Hilbert space (RKHS) [27, 9]. Under this condition, MMD induces a proper metric on distributions, ensuring that any difference between them is detectable [8, 10].
@@
Finally, although $\widehat{\mathrm{MMD}}^{2}$ enjoys favorable theoretical properties, its computation relies on the explicit construction of kernel matrices, resulting in a computational complexity of $\mathcal{O}(d(n+m)^{2})$. Letting $N=n+m$, this complexity can be written as $\mathcal{O}(dN^{2})$ [5]. In large-scale datasets or scenarios where moderate sample sizes require repeated evaluations, such quadratic computational cost severely limits the practical applicability of MMD-based statistical inference.
@@ 2.2 U-statistics and Hoeffding Decomposition
Depending on the sample source, U-statistics can be classified into single-sample U-statistics and generalized U-statistics. Both types form the theoretical foundation of the MMD estimator [26]. Let finite samples be $\mathbf{X}=\{x_{i}\}_{i=1}^{n}$ i.i.d. from $\boldsymbol{X}\sim\mathbb{P}$ and $\mathbf{Y}=\{y_{j}\}_{j=1}^{m}$ i.i.d. from $\boldsymbol{Y}\sim\mathbb{Q}$, where $\boldsymbol{X}$ and $\boldsymbol{Y}$ are independent. Consider a function $\theta(\boldsymbol{X})=\mathbb{E}_{\boldsymbol{X}}[h(x_{1},\dots,x_{r})]$, whose unbiased estimator is
@@
where $r$ denotes the order of the U-statistic and $n$ the sample size. For the single-sample terms in the $\widehat{\mathrm{MMD}}^{2}$ estimator (Eq. 2), the kernel $k(x_{i},x_{i^{\prime}})$ acts on pairs of samples, corresponding to the case where $r=2$.
@@
The orthogonal projection components $g_{c}$ are recursively defined by [14, 4]:
@@
Define $S_{c,n}=\sum_{1\leq i_{1}<\cdots<i_{c}\leq n}g_{c}(x_{i_{1}},\dots,x_{i_{c}})$, then the Hoeffding decomposition yields by [15]:
@@
where $r_{1}$ and $r_{2}$ denote the orders of the U-statistic for the two samples, respectively. For the cross-sample interaction term in the $\widehat{\mathrm{MMD}}^{2}$ estimator, the kernel $k(x_{i},y_{j})$ involves a single observation from each distribution, which corresponds to the case where $r_{1}=1$ and $r_{2}=1$.
@@
Define the partial sums as $S_{c_{1},c_{2},n,m}=\sum_{1\leq i_{1}<\dots<i_{c_{1}}\leq n}\sum_{1\leq j_{1}<\dots<j_{c_{2}}\leq m}g_{c_{1},c_{2}}(x_{i_{1}},\dots,x_{i_{c_{1}}};y_{j_{1}},\dots,y_{j_{c_{2}}})$, the generalized Hoeffding decomposition is then expressed as:
@@ 2.3 Lossless acceleration of the unbiased MMD estimator with the Laplace kernel
Algorithm 1 TriSSL: Triangular Sorted Sum of Laplacians
@@ 3.1 Variance derivation of the unbiased MMD estimator
According to the Hoeffding decomposition theorem (Eq. 8 and 12), the deviation of each component can be expanded into a sum of linear terms and second-order residual terms.
@@
Unlike the approach in [30], which directly applies the general properties of generalized U-statistics by treating the entire MMD as a single monolithic entity, we leverage the natural form of the MMD estimator as a combination of distinct U-statistic components. By combining Eq. 18, 19, and 20, the deviation of the unbiased estimator can be expressed as:
@@
Since different orders in the Hoeffding decomposition are orthogonal, i.e., $\mathrm{Cov}(H_{1},H_{2})=0$ [15], the variance of $\widehat{\mathrm{MMD}}^{2}$ admits the exact decomposition
@@ 3.2.1 Simplification and unbiased estimation of first-order projection terms
where $\widehat{\operatorname{Var}}(\mathcal{U})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$ represent the variances of population projections. To realize this empirically, we replace these population expectations with their empirical projections:
@@
where $\bar{\hat{\mathcal{U}}}=\sum_{i=1}^{n}\hat{\mathcal{U}}(x_{i})\big/n$ and an analogous estimator can be obtained for $\hat{\mathcal{V}}(y_{j})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$:
$\hat{\mathcal{V}}(y_{j})=\frac{1}{m-1}\sum_{j^{\prime}\neq j}k(y_{j},y_{j^{\prime}})-\frac{1}{n}\sum_{i=1}^{n}k(x_{i},y_{j})$ (30)
@@
where we avoid the appearance of complex cross-covariance terms that are prone to estimation bias in finite samples (as seen in [30]). By projecting the estimator onto orthogonal Hoeffding components, we decompose the variance into a streamlined structure of marginal variances and interaction residuals. This not only simplifies the analytical form but also ensures the robustness of the estimator in imbalanced scenarios where $n\neq m$.
@@ 3.2.2 Simplification and unbiased estimation of second-order residual terms
By applying the inclusion-exclusion principle, the unbiased estimator can be expressed entirely in terms of unconstrained matrix operations [28]. Specifically, using the falling factorial $(n)_{k}$ and the centered kernel matrices $\tilde{\mathbf{K}}_{XX}$ and $\tilde{\mathbf{K}}_{YY}$ (with zero diagonals), the expectation $\mathbb{E}[g_{2}^{2}]$ can be written as a linear combination of the squared Frobenius norms, the squared row and column sums, and the squared total sum of the kernel matrices.
@@ 3.2.3 The unbiased estimator of MMD variance
Using the simplification results provided in Eq. 29314344, and 45, the full analytical expression of $\operatorname{Var}(\widehat{\mathrm{MMD}}^{2})$ is given by
@@
Unlike previous works that rely on large-sample asymptotic approximations [30], our derived estimator is exactly unbiased for any finite sample sizes $n$ and $m$. By employing a combinatorial approach based on falling factorials $(n)_{k}$, we explicitly cancel out the systematic bias that typically arises when sample means are substituted for population expectations. This makes our estimator particularly robust in small-sample regimes or highly imbalanced settings ($n\ll m$).
@@
Under the null hypothesis $\mathcal{H}_{0}:\mathbb{P}=\mathbb{Q}$, the unbiased MMD statistic exhibits a "degenerate" behavior. In this regime, the population projections vanish, i.e.,
@@
$\mathcal{V}\left(y\right)$ $:=\mathbb{E}_{\boldsymbol{Y}^{\prime}}\left[k\left(y,\boldsymbol{Y}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{X}}\left[k\left(\boldsymbol{X},y\right)\right]\equiv 0,$ (48)
causing the $O(n^{-1})$ first-order variance $\operatorname{Var}(\mathcal{T}_{1})$ to drop to zero. Consequently, the total population variance is determined solely by the $O(n^{-2})$ second-order residual term $\mathcal{T}_{2}$ [11, 30].
Crucially, this theoretical degeneracy does not imply that the empirical variance of $\hat{\mathcal{U}}(x_{i})$ and $\hat{\mathcal{V}}(y_{j})$ disappears. In finite samples, the empirical projections (Eqs. 28 and 30) retain fluctuations due to the inherent coupling of sample pairs. Our estimator in Eq. (3.2.3) incorporates these fluctuations by integrating $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ into a single analytical form. As a result, our variance estimate is generally larger than those based on asymptotic approximations (e.g., [11]), as it explicitly captures high-order coupling terms that are typically neglected. This "unified" property is a significant advantage: under $\mathcal{H}_{1}$, the $O(n^{-1})$ terms naturally dominate, while under $\mathcal{H}_{0}$, the $O(n^{-2})$ terms automatically take over, which provides a more conservative and robust variance estimate.
@@ 4.2 Accelerated computation of variance components
4.2.1 Accelerated computation of first-order components
@@
4.2.2 Accelerated computation of second-order components
To achieve $O(N)$ computation of the second-order terms, we leverage the bandwidth scaling property (Lemma 4.1). Specifically, while the row sums of the kernel matrix $\mathbf{K}$ are computed using the base bandwidth $\sigma$, the squared kernel terms required for the Frobenius norm $\|\mathbf{K}\|_{F}^{2}=\sum_{i,j}k(x_{i},x_{j})^{2}$ are equivalent to a Laplace kernel with bandwidth $\sigma/2$.
Therefore, we introduce a second pass of prefix-suffix accumulators with $\sigma/2$, denoted as $\{R^{\text{sq}},L^{\text{sq}}\}$ for within-sample matrices and $\{\mathcal{A}^{\text{sq}},\mathcal{Z}^{\text{sq}}\}$ for cross-sample matrices: